OpenStudy (maxdrizner):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3)

OpenStudy (anonymous):

y = 2x + 2 (The slope in y = mx + b is in the m position, therefore the slope of this line is 2) Parallel lines have the same slope. Now we have a choice on how to solve this....there are 2 ways. (1) use y = mx + b (sub in all we know and solve for b) slope(m) = 2 (5,3) x = 5 and y = 3 Now sub.. y = mx + b 3 = 2(5) + b 3 = 10 + b 3 - 10 = b -7 = b Therefore, the parallel line is : y = 2x - 7 (2) This can also be solved by using the point slope formula: y - y1 = m(x - x1) slope(m) = 2 (5,3) x1 = 5 and y1 = 3 now we sub.. y - y1 = m(x - x1) y - 3 = 2(x - 5) y - 3 = 2x - 10 y = 2x - 10 + 3 y = 2x - 7 either way, you will arrive at the correct answer.

OpenStudy (maxdrizner):

thank you

OpenStudy (anonymous):

no problem :) Do you have any questions ?